OpenStudy (anonymous):

solve using substitution method or elemination? 2x+y=-1 -4x+6y=6

OpenStudy (anonymous):

2x + y = -1 -4x + 6y = 6 2 (2x + y = -1) = 4x + 2y = -2 4x + 2y = -2 -4x + 6y = 6 8y = 4 y = 1/2 2x + 1/2 = -1 2x = -1.5 x = -1.5/2 x = -3/4 -4x + 6(1/2) = 6 -4x + 3 = 6 -4x = 3 x = -3/4